OpenStudy (anonymous):

The volume of a cone is 300 cubic kilometers. If the radius of the base is 5 kilometers, what is the height of the cone? Use 3.14 for pi. Show your work and round to the nearest tenth Can someone explain to me how to do this and help me out !?

OpenStudy (anonymous):

V(cone) = 1/3 x pi x r^3 x h

OpenStudy (anonymous):

what does cone mean ?

OpenStudy (anonymous):

The shape that you have ice cream in

OpenStudy (anonymous):

i know but i still dont know how to find the answer :/

OpenStudy (anonymous):

see ive given you the formula you have the values for volume, pi, radius So find the height

OpenStudy (anonymous):

and how do i find the height ?

OpenStudy (anonymous):

3xV/(pi x r^3) = h

OpenStudy (anonymous):

what does v stand for ? and also / is divide , correct ?

OpenStudy (anonymous):

v is the volume

OpenStudy (anonymous):

oh thanks ... one second ...

OpenStudy (anonymous):

i got 2.29 and a bunch of other numbers with the calculator .. is " / " divide ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ohk so is 2.29 my height ?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

ok so what does cone stand for ?

OpenStudy (anonymous):

A cone is type f shape...wiki it

OpenStudy (anonymous):

i need to complete this now , i dont have time to wiki it ....

OpenStudy (anonymous):

What more can i tell you...a cone is a type of solid

OpenStudy (anonymous):

Which has a circular base and tapers into a point at the top

OpenStudy (anonymous):

a cone but in the formula you said V(cone) = 1/3 x pi x r^3 x h what does cone mean in parenthesis ? i know what a cone is ..

OpenStudy (anonymous):

It means Volume of the cone

OpenStudy (anonymous):

so i put 300 there ? or what ...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what would be my final answer ?

OpenStudy (anonymous):

@him1618 ?